Flexible Overhead Budget

Leno Manufacturing Company prepared the following factory overhead cost budget for the Press Department for October of the current year, during which it expected to require 14,000 hours of productive capacity in the department:

Variable overhead cost:
   Indirect factory labor $109,200
   Power and light 4,480
   Indirect materials 44,800
      Total variable overhead cost $158,480
Fixed overhead cost:
   Supervisory salaries $55,470
   Depreciation of plant and equipment 34,870
   Insurance and property taxes 22,190
      Total fixed overhead cost 112,530
Total factory overhead cost $271,010

Assuming that the estimated costs for November are the same as for October, prepare a flexible factory overhead cost budget for the Press Department for November for 12,000, 14,000, and 16,000 hours of production. Round your interim computations to the nearest cent, if required. Enter all amounts as positive numbers.

Leno Manufacturing Company
Factory Overhead Cost Budget-Press Department
For the Month Ended November 30
Direct labor hours 12,000 14,000 16,000
Variable overhead cost:
Indirect factory labor $ $ $
Power and light
Indirect materials
Total variable factory overhead $ $ $
Fixed factory overhead cost:
Supervisory salaries $ $ $
Depreciation of plant and equipment
Insurance and property taxes
Total fixed factory overhead $ $ $
Total factory overhead cost $ $ $
Answer #1
Leno manufacturing company
Factory overhead cost budget - press deptt
Direct labor hours 12000 14000 16000
Variable overhead cost
Indirect factory labor 109200/14000=7.8 93600 109200 124800
power & light 4480/14000 =0.32 3840 4480 5120
indirect material 44800/14000= 3.2 38400 44800 51200
total variable factory overhead (A) 135840 158480 181120
Fixed factory overhead cost
supervisory salaries 55470 55470 55470
Deperication of palnt & equipment 34870 34870 34870
insurance & property taxes 22190 22190 22190
total fixed factory overhead (B) 112530 112530 112530
total factory overhead cost (A)+(B) 248370 271010 293650

Give a thumbs up if it is helpful

please let me know if any doubt